Mathematics 241 – Multivariable Calculus Quiz 4 – October 4, 2013

Your Name:

A) (10) Is α(t) = (cos(t), sin(t)) a flow line of the vector field F (x, y) = (−y, x)? Why or why not?

B) (10) Plot the contours of f(x, y) = x^{2} − y^{2} for the contour values c = −1, 0, 1. Label your contours to show which is which.